If a body moving with simple harmonic motion has an angular velocity of 50 radians per second and an amplitude of 10cm, calculate its linear velocity.
1
Expert's answer
2021-10-13T09:07:44-0400

The displacement of the body in SHM is given in general as follows:


"x(t) = A\sin(\omega t)"

where "A = 10cm" is the amplitude, and "\omega = 50rad/s" is the angular velocity and "t" is time. The linear velocity, by definition, is the derivative from the displacement:


"v(t) = \dfrac{d}{dt}(x(t)) = \dfrac{d}{dt}(A\sin(\omega t)) = A\omega \cos(\omega t)"

Answer. The linear velocity at any given time is: "v(t) = A\omega \cos(\omega t)", where "A = 10cm" is the amplitude, and "\omega = 50rad/s" is the angular velocity and "t" is time.